# Deep Learning-based Intrusion Detection Systems: A Survey

### v2

@@ 5.3. Attack Investigation
In the past, attack investigation is conducted by forward analysis and backward analysis (Hossain et al., 2017). Forward analysis discovers the influence that nodes of interest will cause and backward analysis traces back how nodes of interest are generated. Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns of attack scenario graphs. Table 5 summarizes the overview of attack investigation approaches. Similar to Section 5.2, we exclude papers (Fang et al., 2022; Gao et al., 2018; Ma et al., 2017; Li et al., 2024a; Milajerdi et al., 2019b; Hassan et al., 2020b; Yang et al., 2020; Kwon et al., 2018; Alhanahnah et al., 2022; Ji et al., 2017; Zhu et al., 2023b; Hossain et al., 2017; Wang et al., 2023) slightly relevant to DL for conciseness.
@@
Graph Liu et al. (Liu et al., 2022a) 2022 struc2vec Graph
Neural Karios (Cheng et al., 2024) 2023 GNN Graph
Network TREC (Lv et al., 2024) 2024 GNN Graph
Orthrus (Jiang et al., 2025) 2025 UniMP Path
@@
Unlike detecting intrusive nodes, attack scenario graphs are complicated and thus are hard to handle by traditional learning methods. ProvDetector (Wang et al., 2020) utilizes doc2vec to learn the embedding representation of paths in the provenance graph. Then a density-based detection is deployed to detect abnormal causal paths in the provenance graph.
@@
To capture causal relations within graphs, GNN is commonly adopted. Liu et al. (Liu et al., 2022a) propose an automated attack detection and investigation method via learning the context semantics of the provenance graph. The provenance graph analyzed by struc2vec captures temporal and causal dependencies of system events. Kairos (Cheng et al., 2024) is a practical intrusion detection and investigation tool based on whole-system provenance. Kairos utilizes GNN to analyze system execution history, so that detects and reconstructs complex APTs. It employs a GNN-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event. TREC (Lv et al., 2024) abstracts APT attack investigation problem as a tactics / techniques recognition problem. TREC trains its model in a few-shot learning manner by adopting a Siamese neural network. Orthurus (Jiang et al., 2025) identifies Quality of Attribution as the key factor contributing to whether or not the industry adopts IDS. It first detects malicious hosts using a GNN encoder and then reconstructs the attack path through dependency analysis.
@@ 7.1.2. Insufficient Amount of Data
The second reason for an insufficient amount of data is the limited dataset representativeness. As observed in Table 6, most of the datasets have no more than 10 attack scenarios, not to mention that each of these attack scenarios has been carefully chosen by their authors. This limited number of attack scenarios suggests that existing datasets are almost impossible to represent the diversified attack methods, as the number of CVE records has already been over 280,000 (Corporation, 2025). Furthermore, the existing datasets such as DAPRA TC (DARPA, 2022) are collected in a specific experimental environment and may not cover other types of normal system behaviors. Unicorn SC (Han et al., 2020) is generated by an idealized simulation of supply chain scenarios, which means many real-world features are prone to be ignored in this dataset. Hence, training DL-IDS on these non-representative datasets could be a disaster for the computer systems that they protect.
Finally, the accessibility of datasets further exacerbates the insufficient data problem. Due to privacy and copyright issues, some datasets may be proprietary or difficult to obtain (Wang et al., 2020, 2023). For example, ProvDetector (Wang et al., 2020) conducted a three-month system evaluation in an enterprise environment with 306 hosts and collected benign provenance data of 23 target programs. Yet this dataset has not been made public, rendering it unavailable to improve other DL-IDS and almost all the assessment settings related to ProvDetector are susceptible to inequity.
@@ 7.2. Pre-training Theories and Techniques
With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024b), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023). In contrast, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 6. We can observe that LLMs developed at full speed beginning in 2019. Their prosperity, however, has not extended to DL-IDS. Until now, the only two DL-IDS that incorporate pre-training techniques, AirTag (Ding et al., 2023b) and MAGIC (Jia et al., 2024), still do not make full use of the potential of LLMs. AirTag pre-trains a BERT model on application logs and detects intrusions in terms of embeddings generated by BERT. MAGIC introduces GraphMAE (Hou et al., 2022), a model architecture derived from Graph Autoencoder (Kipf and Welling, 2016) in 2016 but integrated with the famous masked self-supervised learning method (He et al., 2022) in 2022, to conduct self-supervised learning on provenance graphs. MAGIC further designs an adapter to apply the pre-trained model in different detection scenarios. Nevertheless, both AirTag and MAGIC can be regarded as preliminary explorations of pre-training techniques. According to the scaling law (Kaplan et al., 2020), the performance of LLMs will steadily improve, as the parameters, data, and computation increase. And the reasoning ability of LLMs will suddenly emerge (Wei et al., 2022), allowing them to chat with humans smoothly. Such advantageous abilities obviously have not been incorporated into DL-IDS.
Nowadays, some researchers (Ali, 2024; G. Lira et al., 2024; Li et al., 2024b) have started to explore the applications of LLMs on DL-IDS. Yet the theories and techniques of such combination remain challenges. In the following, we will illustrate the identified issues and then point out the future directions.
@@ 7.2.1. Trade-off between Reliability and Generalizability
The governing concern for the employment of LLMs in DL-IDS is reliability (or explainability). Although offering generalizability, LLMs have long been denounced to have issues with hallucinations (Martino et al., 2023; Yao et al., 2023), privacy (Yao et al., 2024) and overreliance (Kim et al., 2024). These unexplainable and uncontrollable features are an absolute disaster for DL-IDS. For example, when feeding log data to LLMs, they sometimes are prone to hallucinate and provide wrong detection results. Attacks thus successfully bypass the detection facilities and can exfiltrate sensitive data in the victim computer systems. Another example for this is that sensitive information may leak from LLMs. Hui et al. (Hui et al., 2024) present a prompt leakage attack for LLMs, which is demonstrated to be effective in both offline settings and real-world LLM applications.
@@ 7.2.2. Short of Statistical Log Modeling
Unfortunately, there are almost no research efforts on statistical modeling of log data, resulting in pre-training techniques of DL-IDS remaining underdeveloped. By contrast, as illustrated in Table 7, the statistical modeling studies of other types of data have already started. Hindle et al. (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language. Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023b, 2024a; Li et al., 2024c, 2022; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection flourish, many of which have already becomes common applications in LLMs. Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
@@ 7.3.1. Limited Forward and Backward Tracing Scope
Thankfully, several pioneering works have studied similar problems (Chen et al., 2024; Manso et al., 2019). In AiVl (Chen et al., 2024), algorithms to bridge log entries and program models are developed using dynamic-static program analysis. Root causes for the exploited vulnerabilities are capable of directly deriving from intrusion detection. Pedro et al. (Manso et al., 2019) investigate detection and mitigation methods for DDoS attacks, aiming to control them immediately. We note that these research attempts are all based on heuristics, either using pre-defined rules to generate root causes, or developing control measures for specific intrusions. Thus, there is a substantial need to introduce advanced DL techniques to this problem.
@@ 7.3.2. Concerns about Data-driven Adversarial Attacks
The robustness of DL-IDS is also challenged by data-driven evasion attacks. To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attacks), making them hard to be detected. By early 2002, David et al. (Wagner and Soto, 2002) have indicated the danger of mimicry attacks on HIDS. Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023) and their studies all present effective evasion of detection. From a study (Chakraborty et al., 2018), DL-IDS can be even plagued by a trivial perturbation in log data. Aware of this issue, R-caid (Goyal et al., 2024) proposes to embed root causes into the detection model for countering adversarial attacks. However, as noted in recent work (Mukherjee et al., 2023; Goyal et al., 2023, 2024), data-driven attacks still remain a major challenge for DL-IDS.
@@ 7.3.3. Underexplored Promising Scenarios
Mobile edge computing (MEC) (Mao et al., 2017; Abbas et al., 2017) is a typical scenario. In the MEC environment, mobile computing, network control, and storage are pushed at the network edges so as to enable computation-intensive tasks at the resource-limited devices. At the network edges, devices such as Unmanned Aerial Vehicles (UAVs) and New Energy Vehicles (NEVs) usually lack computing power and security facilities, making it difficult to prevent them from intrusions (Shrestha et al., 2021). In the meantime, containerized deployment has become one of the dominant ways to deploy microservices. Detecting intrusions on containers is thus of great importance, for which ReplicaWatcher (El Khairi et al., 2024) is a representative work with a special design for microservices. Additionally, industrial networks are characterized by high fidelity, stability, and real-time responsiveness (Knapp, 2024), leading to challenges in adapting DL-IDS to their infrastructures.

### v3

@@ 5.3. Attack Investigation
In the past, attack investigation is conducted by forward analysis and backward analysis (Hossain et al., 2017). Forward analysis discovers the influence that nodes of interest will cause and backward analysis traces back how nodes of interest are generated. Benefiting from DL techniques, both forward and backward analysis can be achieved by learning patterns of attack scenario graphs. Furthermore, visual analytics techniques have been widely used to assist security analysts in understanding the causal chain of intrusions (Zhang et al., 2017; Zhao et al., 2013). Table 5 summarizes the overview of attack investigation approaches. Similar to Section 5.2, we exclude papers (Fang et al., 2022; Gao et al., 2018; Ma et al., 2017; Li et al., 2024a; Milajerdi et al., 2019b; Hassan et al., 2020b; Yang et al., 2020; Kwon et al., 2018; Alhanahnah et al., 2022; Ji et al., 2017; Zhu et al., 2023b; Hossain et al., 2017; Wang et al., 2023) slightly relevant to DL for conciseness.
@@
Liu et al. (Liu et al., 2022a) 2022 struc2vec Graph
Graph Karios (Cheng et al., 2024) 2023 GNN Graph
Neural TREC (Lv et al., 2024) 2024 GNN Graph
Network Orthrus (Jiang et al., 2025) 2025 UniMP Path
Slot (Qiao et al., 2025) 2025 GNN Graph
FeCoGraph (Mao et al., 2025) 2025 GCN Graph
@@
Unlike detecting intrusive nodes, attack scenario graphs are complicated and thus are hard to handle by traditional learning methods. ProvDetector (Wang et al., 2020) utilizes doc2vec to learn the embedding representation of paths in the provenance graph. Then a density-based detection is deployed to detect abnormal causal paths in the provenance graph. BehaviorBaseline (Zhu et al., 2025) presents a novel learning-based anomaly detection method for large-scale provenance graphs. It incorporates dynamic graph processing with adaptive encoding and a tag-propagation framework for real-time detction.
@@
To capture causal relations within graphs, GNN is commonly adopted. Liu et al. (Liu et al., 2022a) propose an automated attack detection and investigation method via learning the context semantics of the provenance graph. The provenance graph analyzed by struc2vec captures temporal and causal dependencies of system events. Kairos (Cheng et al., 2024) is a practical intrusion detection and investigation tool based on whole-system provenance. Kairos utilizes GNN to analyze system execution history, so that detects and reconstructs complex APTs. It employs a GNN-based encoder-decoder architecture to learn the temporal evolution of provenance graph structure changes and quantify the abnormal degree of each system event. TREC (Lv et al., 2024) abstracts APT attack investigation problem as a tactics / techniques recognition problem. TREC trains its model in a few-shot learning manner by adopting a Siamese neural network. Orthurus (Jiang et al., 2025) identifies Quality of Attribution as the key factor contributing to whether or not the industry adopts IDS. It first detects malicious hosts using a GNN encoder and then reconstructs the attack path through dependency analysis. Slot (Qiao et al., 2025), based on provenance graphs and graph reinforcement learning, uncovers hidden relationships among system behaviors, dynamically adapts to new activities and attack strategies, resists adversarial attacks, and automatically constructs attack chains. FeCoGraph (Mao et al., 2025) directly processes traffic embedding through line graphs to adapt to various GNNs, covering more attack scenarios while protecting data privacy.
@@ 6.1. Dimensions of Datasets
  • Operating Systems: The operating system determines the generalizability of the dataset. The more operating systems a dataset covers and the more common they are, the more comprehensively it can evaluate PIDS performance.
@@ 6.2. Public Datasets
ATLAS dataset (Alsaheel et al., 2021) implements 10 attacks based on their detailed reports on real-world APT campaigns and generates audit logs in a controlled testbed environment. Among the ten attacks, four are from single host and the rest six are from multiple hosts. All attacks were developed and executed on Windows 7 32-bit virtual machines and took an hour to complete, along with a 24-hour-window audit logs for benign system behaviors.
ATLASv2 dataset (Riddle et al., 2023) enriches the ATLAS dataset with higher quality background noise and additional logging vantage points. In this dataset, two researchers use the victim machines as their primary work stations throughout the course of engagement, instead of depending on automated scripts to generate activity. System logging, in contrast, cover a five-day period, where the first four days simulate normal work days and the fifth day begins with benign activity then trasitions into execution of the corresponding attack.
@@
AutoLabel dataset (Peng et al., 2025) automates fine-grained log labeling by reducing the labeling problem to obtaining an accurate attack subgraph in a provenance graph. Its experiments consist of 29 scenarios, including 25 real CVE vulnerabilities across 12 widely-used applications (spanning 5 programming languages) plus a Sandworm threat simulation by MITRE CTID.
@@ 7.1.2. Insufficient Amount of Data
The second reason for an insufficient amount of data is the limited dataset representativeness. As observed in Table 6, most of the datasets have no more than 10 attack scenarios, not to mention that each of these attack scenarios has been carefully chosen by their authors. This limited number of attack scenarios suggests that existing datasets are almost impossible to represent the diversified attack methods, as the number of CVE records has already been over 280,000 (Corporation, 2025). Furthermore, the existing datasets such as DAPRA TC E3 (DARPA, 2022) are collected in a specific experimental environment and may not cover other types of normal system behaviors, and are proven that a significant amount of synthetic data exists (Liu et al., 2025a). DARPA TC E5 (DARPA, 2022) is unusable for most experiments due to the sparse and error-filled documentation. Unicorn SC (Han et al., 2020) is generated by an idealized simulation of supply chain scenarios, which means many real-world features are prone to be ignored in this dataset. Hence, training DL-IDS on these non-representative datasets could be a disaster for the computer systems that they protect.
Finally, the accessibility of datasets further exacerbates the insufficient data problem. Due to privacy and copyright issues, some datasets may be proprietary or difficult to obtain (Wang et al., 2020, 2023). Moreover, ProvDetector (Wang et al., 2020) conducted a three-month system evaluation in an enterprise environment with 306 hosts and collected benign provenance data of 23 target programs. Yet this dataset has not been made public, rendering it unavailable to improve other DL-IDS and almost all the assessment settings related to ProvDetector are susceptible to inequity.
@@ 7.2. Pre-training Theories and Techniques
With the development of pre-training techniques, LLMs have been adopted in many fields such as finance (Zhao et al., 2024c), education (Neumann et al., 2024), medicine (Peng et al., 2023), and even other domains of cybersecurity (Cummins et al., 2025; Hu et al., 2024; Gu, 2023). In contrast, the adoption of LLMs in DL-IDS is stagnant, as shown in Figure 6. We can observe that LLMs developed at full speed beginning in 2019. Their prosperity, however, has not extended to DL-IDS. Until now, the only two DL-IDS that incorporate pre-training techniques, AirTag (Ding et al., 2023b) and MAGIC (Jia et al., 2024), still do not make full use of the potential of LLMs. AirTag pre-trains a BERT model on application logs and detects intrusions in terms of embeddings generated by BERT. MAGIC introduces GraphMAE (Hou et al., 2022), a model architecture derived from Graph Autoencoder (Kipf and Welling, 2016) in 2016 but integrated with the famous masked self-supervised learning method (He et al., 2022) in 2022, to conduct self-supervised learning on provenance graphs. MAGIC further designs an adapter to apply the pre-trained model in different detection scenarios. Nevertheless, both AirTag and MAGIC can be regarded as preliminary explorations of pre-training techniques. According to the scaling law (Kaplan et al., 2020), the performance of LLMs will steadily improve, as the parameters, data, and computation increase. And the reasoning ability of LLMs will suddenly emerge (Wei et al., 2022), allowing them to chat with humans smoothly. Such advantageous abilities obviously have not been incorporated into DL-IDS.
Nowadays, some researchers (Ali, 2024; G. Lira et al., 2024; Li et al., 2024b; Mukherjee and Kantarcioglu, 2025) have started to explore the applications of LLMs on DL-IDS. Yet the theories and techniques of such combination remain challenges. In the following, we will illustrate the identified issues and then point out the future directions.
@@ 7.2.1. Trade-off between Reliability and Generalizability
The governing concern for the employment of LLMs in DL-IDS is reliability (or explainability). Although offering generalizability, LLMs have long been denounced to have issues with hallucinations (Martino et al., 2023; Yao et al., 2023), privacy (Yao et al., 2024; Yang et al., 2025; He et al., 2025), overreliance (Kim et al., 2024), and backdoor threats (Liu et al., 2025b). These unexplainable and uncontrollable features are an absolute disaster for DL-IDS. For example, when feeding log data to LLMs, they sometimes are prone to hallucinate and provide wrong detection results. Attacks thus successfully bypass the detection facilities and can exfiltrate sensitive data in the victim computer systems. Another example for this is that sensitive information may leak from LLMs. Hui et al. (Hui et al., 2024) present a prompt leakage attack for LLMs, which is demonstrated to be effective in both offline settings and real-world LLM applications.
@@ 7.2.2. Short of Statistical Log Modeling
Unfortunately, there are almost no research efforts on statistical modeling of log data, resulting in pre-training techniques of DL-IDS remaining underdeveloped. By contrast, as illustrated in Table 7, the statistical modeling studies of other types of data have already started. Hindle et al. (Hindle et al., 2016) demonstrate that the source code is very repetitive and predictable, and, in fact, even more so than natural language. Driven by such statistical modeling conclusion, DL-based source code applications (Xu et al., 2023b, 2024a; Li et al., 2024c, 2022a; Svyatkovskiy et al., 2020; Feng et al., 2020; Gu et al., 2025) such as code generation and code clone detection flourish, many of which have already becomes common applications in LLMs. Similar cases can be found for speech data, whose applications are like text to speech (Ren et al., 2019; Guo et al., 2024; Park et al., 2025) and speech recognition (Baevski et al., 2021).
@@ 7.3.1. Limited Forward and Backward Tracing Scope
Thankfully, several pioneering works have studied similar problems (Chen et al., 2024; Manso et al., 2019). In AiVl (Chen et al., 2024), algorithms to bridge log entries and program models are developed using dynamic-static program analysis. Root causes for the exploited vulnerabilities are capable of directly deriving from intrusion detection. Pedro et al. (Manso et al., 2019) investigate detection and mitigation methods for DDoS attacks, aiming to control them immediately. Additionally, semi-automated adaptive network defense (SAND) (Chen et al., 2022) leverages SDN to dynamically generate and deploy defense rules. We note that these research attempts are all based on heuristics, either using pre-defined rules to generate root causes, or developing control measures for specific intrusions. Thus, there is a substantial need to introduce advanced DL techniques to this problem.
@@ 7.3.2. Concerns about Data-driven Adversarial Attacks
The robustness of DL-IDS is also challenged by data-driven evasion attacks. To evade the detection, the malicious behaviors usually mimic the benign ones (a.k.a., mimicry attacks), making them hard to be detected. By early 2002, David et al. (Wagner and Soto, 2002) have indicated the danger of mimicry attacks on HIDS. Recently, researchers have started to investigate mimicry attacks on DL-IDS (Mukherjee et al., 2023; Goyal et al., 2023; Liu et al., 2022c) and their studies all present effective evasion of detection. From a study (Chakraborty et al., 2018), DL-IDS can be even plagued by a trivial perturbation in log data. Aware of this issue, R-caid (Goyal et al., 2024) proposes to embed root causes into the detection model for countering adversarial attacks. However, as noted in recent work (Mukherjee et al., 2023; Goyal et al., 2023, 2024), data-driven attacks still remain a major challenge for DL-IDS.
@@ 7.3.3. Underexplored Promising Scenarios
Mobile edge computing (MEC) (Mao et al., 2017; Abbas et al., 2017; Li et al., 2022b) is a typical scenario. In the MEC environment, mobile computing, network control, and storage are pushed at the network edges so as to enable computation-intensive tasks at the resource-limited devices. At the network edges, devices such as Unmanned Aerial Vehicles (UAVs) and New Energy Vehicles (NEVs) usually lack computing power and security facilities, making it difficult to prevent them from intrusions (Shrestha et al., 2021). In the meantime, containerized deployment has become one of the dominant ways to deploy microservices. Detecting intrusions on containers is thus of great importance, for which ReplicaWatcher (El Khairi et al., 2024) is a representative work with a special design for microservices. Additionally, industrial networks are characterized by high fidelity, stability, and real-time responsiveness (Knapp, 2024), leading to challenges in adapting DL-IDS to their infrastructures.